# Closed-Form Approximation of the Total Variation Proximal Operator

### v1

@@ 4.2 Computed Tomography
We validate the application of ${\mathcal{S}}_{\tau}({\bm{z}})$ for the limited-angle computed tomography (CT) reconstruction. The data fidelity term in Eq. (2) is $g({\bm{x}})=\frac{1}{2}\|{\bm{A}}{\bm{x}}-{\bm{y}}\|_{2}^{2},$ where ${\bm{A}}$ represents the CT imaging forward operator using only $45$ angles, resulting in an ill-posed inverse problem. We show the experimental results for the regularization parameters $\lambda$ corresponding to $2.5$, $5$, and $10$, using both APGM and ADMM. ADMM is implemented to solve the constrained optimization problem
@@ 4.3 Discussion
According to Proposition 1, ${\mathcal{S}}_{\tau}({\bm{z}})$ serves as the proximal operator for some convex function. Consequently, the iterates of a proximal-based reconstruction algorithm a convex function are expected to converge. Activation of the termination criteria in 9 occurs for all the experiments, validating the convergence of the algorithms under ${\mathcal{S}}_{\tau}({\bm{z}})$.
The statements in Proposition 2 and Corollary 1 imply smaller $\tau$ in ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation with respect to the true TV reconstruction. This relationship is experimentally validated in Table 1 across all experiments: a smaller $\tau$ results in a smaller relative error in the approximate reconstruction loss $\widehat{f}$, compared to the true TV reconstruction loss $f^{*}$. Similarly, the PSNR with respect to the true TV reconstruction increases as $\tau$ decreases, suggesting ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation of the TV proximal operator. Figure 2 illustrates the dependence of cost accuracy on parameter $\gamma$ for image denoising with $\lambda=0.5$ and CT image reconstruction with $\lambda=5$. Figures 1,  3, and 4 illustrate visual comparison for image denoising and CT reconstruction using true TV proximal and ${\mathcal{S}}_{\tau}({\bm{z}})$ operator with various $\gamma$. The visual results indicate that ${\mathcal{S}}_{\tau}({\bm{z}})$ promotes piecewise smooth images, as is expected from an approximate TV proximal operator. Both visual and numerical results from the experiments support the theoretical finding presented in Section 3.2, demonstrating that ${\mathcal{S}}_{\tau}({\bm{z}})$ can approximate the TV proximal and be effectively integrated into proximal-based iterative algorithms to achieve approximate TV-regularized reconstructions.
@@ 6 Appendix
In this section, we will prove Propositions  12, and Corollary 1. First, we define the function $\bar{h}:\mathbb{R}^{2nd}\rightarrow\mathbb{R}$, where $\bar{h}$ has an anisotropic and isotropic form, $\bar{h}^{\text{a}}$ and $\bar{h}^{\text{i}}$, respectively. For ${\bm{u}}\in\mathbb{R}^{2nd}$,
$\bar{h}^{\text{a}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{1,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}$ (10)
$\bar{h}^{\text{i}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{2,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{2}\;,$ (11)
where ${\bm{u}}^{\text{dif}}\in\mathbb{R}^{nd}$ denotes the difference coefficients. $\|{\bm{u}}^{\text{dif}}\|_{p,1}$ is a group norm, where $p$ is either $1$ or $2$, corresponding to anisotropic and isotropic TV, respectively. The grouping is based on the $d$ elements that represent the discrete differences at a specific location for each of the $d$ dimensions, with the $p$-norm applied within each group. Each $[{\bm{u}}^{\text{dif}}]_{i}\in\mathbb{R}^{d}$ is the vector of discrete differences at location $i$. We use
@@
holds. Note that Eq. (14) implies that for ${\bm{z}}\in\mathbb{R}^{n}$, $\bar{h}$ evaluated at ${\bm{W}}{\bm{z}}$ is equivalent to ${h}$ evaluated at ${\bm{z}}$; however, $\bar{h}$ is not the TV function itself since it is defined on the entire $\mathbb{R}^{2nd}$, and not $\mathbb{R}^{n}$.
@@
[…]
@@
Next, letting $\psi({\bm{z}})=[\phi\circ{\bm{W}}]({\bm{z}})$, we need to show $\psi$ is closed and convex. Since $\phi$ is closed, $\phi\circ{\bm{W}}$ is also closed. To show convexity, letting ${\bm{z}}_{1},{\bm{z}}_{2}\in\mathbb{R}^{n}$,
$\psi(\alpha{\bm{z}}_{1}+(1-\alpha){\bm{z}}_{2})$ $=[\phi\circ{\bm{W}}](\alpha{\bm{z}}_{1}+(1-\alpha){\bm{z}}_{2})$
$=\phi(\alpha{\bm{W}}{\bm{z}}_{1}+(1-\alpha){\bm{W}}{\bm{z}}_{2})$
$\leq\alpha\phi({\bm{W}}{\bm{z}}_{1})+(1-\alpha)\phi({\bm{W}}{\bm{z}}_{2})$
$=\alpha\psi({\bm{z}}_{1})+(1-\alpha)\psi({\bm{z}}_{2})\;,$ (16)
where the inequality comes from convexity of $\phi$. By definition of convexity, $\psi$ is thus convex. We can show $\|{\bm{W}}\|_{2}=1$ as follows
@@
[…]
$\quad=\|{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
@@
Proof of Part (a). By using the definition of TV in Eq. (13) we obtain
${h}({\mathcal{S}}_{\tau}({\bm{z}}))$ $=\|{\bm{D}}\hskip 1.49994pt{\mathcal{S}}_{\tau}({\bm{z}})\|_{p,1}$
[…]
$=2\sqrt{d}\hskip 1.49994pt\|[{\bm{W}}({\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}))]^{\text{dif}}\|_{p,1}\;,$
where we used the definition of the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ in the second line. In the last line, we used the fact that ${\bm{W}}$ includes difference operator ${\bm{D}}$ scaled by $2\sqrt{d}$ (that is, ${\bm{W}}\,\coloneqq\,(1/2\sqrt{d})[{\bm{A}}\quad{\bm{D}}]^{\mathsf{T}}$).
The thresholding operator ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}$ reduces the magnitude of each difference component by multiplying it by some value in $(0,1)$. Each average component remains the same. Thus, $[{\bm{W}}{\bm{z}}]_{i}$ is multiplied by $\alpha_{i}$, where $0<\alpha_{i}\leq 1$. Therefore, we can replace ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ with the vector ${\bm{\alpha}}\in\mathbb{R}^{2nd}$ where each $\alpha_{i}$ satisfies $0\leq\alpha_{i}\leq 1$.Here, $\odot$ signifying component-wise multiplication. Consequently, we have
${h}({\mathcal{S}}_{\tau}({\bm{z}}))$ $=2\sqrt{d}\hskip 1.49994pt\|[{\bm{W}}{\bm{W}}^{{\mathsf{T}}}({\bm{\alpha}}\odot({\bm{W}}{\bm{z}}))]^{\text{dif}}\|_{p,1}$
$=2\sqrt{d}\hskip 1.49994pt\|[{\bm{\alpha}}]^{\text{dif}}\odot[{\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{W}}{\bm{z}}]^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\hskip 1.49994pt\|[{\bm{W}}{\bm{z}}]^{\text{dif}}\|_{p,1}$
$=\|{\bm{D}}{\bm{z}}\|_{p,1}={h}({\bm{z}})\;.$
Component-wise multiplication is commutative, allowing us to swap ${\bm{\alpha}}$ and ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ in the second equality. The inequality comes from ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}={\bm{I}}$ and the fact that each $\alpha_{i}$ satisfies $0\leq\alpha_{i}\leq 1$. This establishes the desired result of part (a).
Proof of Part (b). We know $\|{\bm{W}}\|_{2}=1$ from (17). For both the anisotropic and isotropic versions, by its definition the soft-thresholding operator ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ at most changes each of the $nd$ difference elements of ${\bm{W}}{\bm{z}}$ by $2\tau\sqrt{d}$, so we can bound
$\|{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}$ $\leq\sqrt{nd(2\tau\sqrt{d})^{2}}$
@@
$\leq\Big{\|}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\Big{\|}_{2}$
@@
$d(\bm{0},\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$
$\quad\leq\Big{\|}\bm{0}-\left(\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\right)\Big{\|}_{2}$
$\quad\leq\frac{1}{\tau}\Big{\|}{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}}\Big{\|}_{2}+\Big{\|}{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\Big{\|}_{2}$
$\quad\leq\frac{1}{\tau}(2\tau d\sqrt{n})+\|{\bm{g}}_{1}\|_{2}$
$\quad\leq\frac{1}{\tau}(2\tau d\sqrt{n})+2d\sqrt{n}=\frac{\tau\epsilon_{1}}{\tau}\;,$ (20)
@@
Proof of Part (c). From the definition of the subgradient and the proximal operator, we have

### v2

@@ 4.2 Computed Tomography
We validate the application of ${\mathcal{S}}_{\tau}({\bm{z}})$ for the limited-angle computed tomography (CT) reconstruction. The data fidelity term in Eq. (2) is $g({\bm{x}})=\frac{1}{2}\|{\bm{A}}{\bm{x}}-{\bm{y}}\|_{2}^{2},$ where $\bm{A}$} represents the CT imaging forward operator using only $45$ angles, resulting in an ill-posed inverse problem. We show the experimental results for the regularization parameters $\lambda$ corresponding to $2.5$, 55, and $10$, using both APGM and ADMM. ADMM is implemented to solve the constrained optimization problem
@@ 4.3 Discussion
According to Proposition 1, ${\mathcal{S}}_{\tau}({\bm{z}})$ serves as the proximal operator for some convex function. Consequently, proximal-based reconstruction algorithms using this operator are guaranteed to converge. Activation of the termination criteria in 13 occurs for all the experiments, validating the convergence of the algorithms under ${\mathcal{S}}_{\tau}({\bm{z}})$.
@@
The statements in Proposition 2 and Proposition 3 imply that a smaller $\tau$ in ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation with respect to the true TV reconstruction. This relationship is experimentally validated in Table 1 across all experiments: a smaller $\tau$ results in a smaller relative error in the approximate reconstruction loss $\widehat{f}$, compared to the true TV reconstruction loss $f^{*}$. Relative error, denoted as Rel. Err. in Table 1, is defined to be
$\frac{\widehat{f}-f^{*}}{f^{*}}.$
Similarly, the PSNR with respect to the true TV reconstruction increases as $\tau$ decreases, suggesting ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation of the TV proximal operator. Table 1 also displays the trade-off between approximation accuracy and acceleration when using ${\mathcal{S}}_{\tau}({\bm{z}})$ compared to $50$ iterations of FPG for the exact proximal. “Accel” is computed as the number of iteration using ${\mathcal{S}}_{\tau}({\bm{z}})$ divided by total number of sub-iterations when using FPG for the exact proximal. Figure 2 illustrates the dependence of cost accuracy on parameter $\gamma$ for image denoising with $\lambda=0.5$ and CT image reconstruction with $\lambda=5$. Figures 1,  3, and 4 illustrate visual comparison for image denoising and CT reconstruction using true TV proximal and ${\mathcal{S}}_{\tau}({\bm{z}})$ operator with various $\gamma$. The visual results indicate that ${\mathcal{S}}_{\tau}({\bm{z}})$ promotes piecewise smooth images, as is expected from an approximate TV proximal operator. Both visual and numerical results from the experiments support the theoretical finding presented in Section 3.2, demonstrating that ${\mathcal{S}}_{\tau}({\bm{z}})$ can approximate the TV proximal and be effectively integrated into proximal-based iterative algorithms to achieve approximate TV-regularized reconstructions.
@@ 5 Conclusion
Acknowledgment
The authors thank the anonymous reviewers for their careful reading and comments that helped improve the final version of this manuscript.
@@ 6 Appendix
In this section, we will prove Propositions  12, and  3. The linear map ${\bm{W}}:R^{n}\rightarrow\mathbb{R}^{2nd}$ is defined as
${\bm{W}}:=\frac{1}{2\sqrt{d}}\begin{bmatrix}{\bm{M}}\vskip 3.0pt\\ {\bm{D}}\end{bmatrix},\qquad\begin{aligned} {\bm{M}}&:=\big{[}{\bm{M}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{M}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\\ {\bm{D}}&:=\big{[}{\bm{D}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{D}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\end{aligned}$
where $\bm{M}$} is the averaging map and $\bm{D}$} is the finite difference map. Therefore, for any vector ${\bm{u}}\in\mathbb{R}^{2nd}$ half of the components correspond to the averaging map and the remaining components correspond to the finite difference map. We use the notation ${\bm{u}}^{\text{dif}}$ to correspond to the difference coefficients.
Using this notation, we define the function $\bar{h}:\mathbb{R}^{2nd}\rightarrow\mathbb{R}$, where $\bar{h}$ has an anisotropic and isotropic form, $\bar{h}^{\text{a}}$ and $\bar{h}^{\text{i}}$, respectively. Let $G_{1},...,G_{n}$ be a partition of $\{1,2,...,nd\}$ such that each $G_{i}$ is the set of indices corresponding to the finite difference of pixel $i$. For ${\bm{u}}\in\mathbb{R}^{2nd}$,
$\bar{h}^{\text{a}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{1,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{G_{i}}\|_{1}$ (14)
$\bar{h}^{\text{i}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{2,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{G_{i}}\|_{2}\;,$ (15)
where $\|{\bm{u}}^{\text{dif}}\|_{p,1}$ is a group norm, with $p$ either $1$ or $2$, to anisotropic and isotropic TV, respectively. The grouping is based on the $d$ elements that represent the discrete differences at a specific location for each of the $d$ dimensions, with the $p$-norm applied within each group. Each ${\bm{u}}^{\text{dif}}]_{G_{i}}\in\mathbb{R}^{d}$ is the vector at the indices given by the set $G_{i}$. We use
@@
holds. Note that Eq. (18) implies that for ${\bm{z}}\in\mathbb{R}^{n}$, $\bar{h}$ evaluated at ${\bm{W}}{\bm{z}}$ is equivalent to h{h} evaluated at $\bm{z}$}; however, $\bar{h}$ is not the TV function itself since it is defined on the entire $\mathbb{R}^{2nd}$, and not $\mathbb{R}^{n}$.
@@
${\mathsf{prox}}_{\tau\bar{h}}({\bm{u}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{u}})\quad\forall{\bm{u}}\in\mathbb{R}^{2nd}\;.$ (19)
@@
Next, letting $\psi({\bm{z}})=[\phi\circ{\bm{W}}]({\bm{z}})$, we need to show $\psi$ is closed and convex. Since $\phi$ is closed, $\phi\circ{\bm{W}}$ is also closed. Since $\phi$ is convex and $\bm{W}$} is linear, $\phi\circ{\bm{W}}$ is convex. We can show $\|{\bm{W}}\|_{2}=1$ as follows
@@
$\quad=\|{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
$\quad=\|{\bm{W}}^{{\mathsf{T}}}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2}))\|_{2}$
$\quad\leq\|{\bm{W}}^{{\mathsf{T}}}\|_{2}\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
$\quad=\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
@@
Let $\widetilde{h}^{a}({\bm{z}})$ and $\widetilde{h}^{i}({\bm{z}})$ denote the Huber variants of anisotropic TV and isotropic TV given in Eq. (9) and Eq. (11), respectively. Then,
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{a}({\bm{z}})\quad\text{and}\quad{\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{i}({\bm{z}}).$
Proof.
We will first consider anisotropic TV ${\mathcal{S}}^{a}_{\tau}({\bm{z}})$. The soft-thresholding operator can be rewritten as
${\mathcal{T}}^{\hskip 1.22911pta}_{\lambda}(u)\,\coloneqq\,{\mathsf{max}}(|u|-\lambda,0)\frac{u}{|u|}=u-\mathop{\mathsf{min}}(\lambda,|u|)\text{sgn}(u).$
Since soft-thresholding only affects the difference components, it can be written as
${\mathcal{S}}_{\tau}({\bm{z}})$ $={\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}^{\hskip 1.22911pta}_{\tau}({\bm{W}}{\bm{z}})$
[…]
$={\bm{z}}-\left[\frac{{\bm{M}}^{{\mathsf{T}}}}{2\sqrt{d}}\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\right]\begin{bmatrix}0\\ \vdots\\ 0\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{1}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{1}}{2\sqrt{d}}\right)\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{nd}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{nd}}{2\sqrt{d}}\right)\end{bmatrix}$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\begin{bmatrix}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{1}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{1}}{2\sqrt{d}}\right)\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{nd}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{nd}}{2\sqrt{d}}\right)\end{bmatrix}$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|{\bm{D}}{\bm{z}}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{{\bm{D}}{\bm{z}}}{2\sqrt{d}}\right)$
$={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\mathop{\mathsf{min}}\left(1,\frac{|{\bm{D}}{\bm{z}}|}{\tau 4d}\right)\text{sgn}\left({\bm{D}}{\bm{z}}\right).$
In the last two lines, the operations $\mathop{\mathsf{min}}$ and sgn are to be understood as being applied element-wise to ${\bm{D}}{\bm{z}}$. The function $\phi_{\theta}$ from Eq. (10) has the derivative,
$\phi_{\theta}^{\prime}(u)=\begin{cases}\frac{u}{\theta}&\text{for }|u|\leq\theta\\ \text{sgn}(u)&\text{for }|u|>\theta\end{cases}=\mathop{\mathsf{min}}\left(1,\frac{|u|}{\theta}\right)\text{sgn}(u)$
Therefore, the operator can be written as
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{\tau 4d}\left({\bm{D}}{\bm{z}}\right),$
where $\phi^{\prime}_{\tau 4d}$ is applied element-wise. In other words, ${\mathcal{S}}^{a}_{\tau}({\bm{z}})$ is a gradient step with step-size $\tau$ on
$\widetilde{h}^{a}({\bm{z}})=\sum_{i=1}^{nd}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{i}).$ (21)
A step-size of $\tau$ satisfies the criterion for the step size to be $\frac{1}{L}$, where $L$ is the Lipschitz constant of the gradients. The Lipschitz constant for Eq. 21 is $1/\tau$, which can be verified as follows:
$\|\nabla\widetilde{h}({\bm{x}})-\nabla\widetilde{h}({\bm{y}})\|$ $=\|{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{x}})-{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{y}})\|$
$\leq\|{\bm{D}}\|\|\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{x}})-\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{y}})\|$
$\leq\left(\frac{1}{4\tau d}\right)\|{\bm{D}}\|\|{\bm{D}}{\bm{x}}-{\bm{D}}{\bm{y}}\|$
$\leq\left(\frac{1}{4\tau d}\right)\|{\bm{D}}\|^{2}\|{\bm{x}}-{\bm{y}}\|$
$=\frac{1}{\tau}\|{\bm{x}}-{\bm{y}}\|,$
where we use the fact that the finite difference matrix $\bm{D}$} has spectral norm $2\sqrt{d}$.
Now consider isotropic ${\mathcal{S}}^{i}_{\tau}({\bm{z}})$. The soft-thresholding operator can be rewritten as
${\mathcal{T}}^{\hskip 1.22911pt\text{i}}_{\lambda}({\bm{u}}):={\mathsf{max}}(\|{\bm{u}}\|_{2}-\lambda,0)\frac{{\bm{u}}}{\|{\bm{u}}\|_{2}}={\bm{u}}-\mathop{\mathsf{min}}(\lambda,\|{\bm{u}}\|)\frac{{\bm{u}}}{\|{\bm{u}}\|}.$
Let $G_{i}$ for $i\in\{1,...,n\}$ be the set of $d$ indices corresponding to the differences along all $d$ dimensions for $x_{i}$. Since soft-thresholding affects only the difference components, it can be rewritten
${\mathcal{S}}_{\tau}({\bm{z}})$ $={\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}^{\hskip 1.22911pti}_{\tau}({\bm{W}}{\bm{z}})$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\begin{bmatrix}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{\|[{\bm{D}}{\bm{z}}]_{G_{1}}\|}{2\sqrt{d}}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{1}}}{\|[{\bm{D}}{\bm{z}}]_{G_{1}}\|}\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{\|[{\bm{D}}{\bm{z}}]_{G_{n}}\|}{2\sqrt{d}}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{n}}}{\|[{\bm{D}}{\bm{z}}]_{G_{n}}\|}\end{bmatrix}$
$={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\mathop{\mathsf{min}}\left(1,\frac{\|[{\bm{D}}{\bm{z}}]_{G_{i}}\|}{\tau 4d}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{i}}}{\|[{\bm{D}}{\bm{z}}]_{G_{i}}\|}.$
The last line is meant to be understood as applying $\mathop{\mathsf{min}}$ and the norm to each grouping of indices. The function $\psi_{\theta}$ from Eq. (12) has the gradient
$\nabla\psi_{\theta}({\bm{u}})=\begin{cases}\frac{{\bm{u}}}{\theta}&\text{for }\|{\bm{u}}\|\leq\theta\\ \frac{{\bm{u}}}{\|{\bm{u}}\|}&\text{for }\|{\bm{u}}\|>\theta\end{cases}=\mathop{\mathsf{min}}\left(1,\frac{\|{\bm{u}}\|}{\theta}\right)\frac{{\bm{u}}}{\|{\bm{u}}\|}$
and so the operator can be written as
${\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\nabla\phi_{\tau 4d}({\bm{D}}{\bm{z}}),$
where $\nabla\phi_{\tau 4d}$ is applied to each grouping of indices. That is, ${\mathcal{S}}^{i}_{\tau}({\bm{z}})$ is a gradient step with the step-size $\tau$ on
$\widetilde{h}^{i}({\bm{z}})=\sum_{i=1}^{n}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{G_{i}}).$ (22)
Again, $\tau$ satisfies the step-size to be $\frac{1}{L}$, since the Lipschitz constant of the gradients for Eq. 22 remains $1/\tau$. ∎
Proof of Proposition 3
Proposition 3.
@@
Proof of Part (a). We know $\|{\bm{W}}\|_{2}=1$ from (20). For both the anisotropic and isotropic versions, by its definition the soft-thresholding operator ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ at most changes each of the $nd$ difference elements of ${\bm{W}}{\bm{z}}$ by $2\tau\sqrt{d}$, so we can bound
$\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}$ $\leq\sqrt{nd(2\tau\sqrt{d})^{2}}$
@@
$\leq\Big{\|}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\Big{\|}_{2}$
@@
$d(\bm{0},\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$ $\leq\Big{\|}\bm{0}-\left(\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\right)\Big{\|}_{2}$
$\leq\frac{1}{\tau}\Big{\|}{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}}\Big{\|}_{2}+\Big{\|}{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\Big{\|}_{2}$
$\leq\frac{1}{\tau}(2\tau d\sqrt{n})+\|{\bm{g}}_{1}\|_{2}$
$\leq\frac{1}{\tau}(2\tau d\sqrt{n})+2d\sqrt{n}=\frac{\tau\epsilon_{1}}{\tau}\;,$ (25)
@@
Proof of Part (b). From the definition of the subgradient and the proximal operator, we have